A loan of $8 000 is to be amortized with equal monthly payments over 2 years at j12 = 15%. Find the outstanding balance after 7 months and split the eighth payment into interest and principal portions.
1
Expert's answer
2020-04-01T09:22:21-0400

The monthly payment is

R = 8000 / a24|.0125 = $387.8931844 "\approx" $387.90

The outstanding principal P after 7 payments is

P = 8000(1.0125)7 - 387.90s7|.0125 = $5907.53

The interest portion of the eight payment is

(5907.53)(0.0125) = $73.84

And the principal portion is 387.90 - 73.84 = $314.06